Ed has 100 dollars more than Robert. After Ed spends 20 dollars on groceries, Ed has 5 times as much as money as Robert. How much money does he have?: Ed has 100 dollars more than Robert. After Ed spends 20 dollars on groceries, Ed has 5 times as much as money as Robert. How much money does he have?
Answer by ptaylor(1332)

Let x=amount that Robert has
Then Ed has x+100
So, our equation to solve is:
(x+100)-20=5x get rid of parens
x+100-20=5x
x+80=5x subtract x from each side
x-x+80=5x-x collect like terms
80=4x divide each side by 4
x=20 dollars-------------------amount that Robert has
x+100=20+100=120 dollars ----amount that Ed has
CK
120-20=5*20
100=100
Hope this helps---ptaylor